OpenStudy (anonymous):

Suppose the coefficient of kinetic firxxtion between mass a and a plane is .10 and mass A and B is 2.7kg. As Mass B moves down, determine the magnitude of mass A and B given an angle of 31 degrees. What is the smallest value of kinetic friction to keep the system from accelerating?

OpenStudy (anonymous):

I've figured out wha the acceleration is (2.0m/s^2 \[(2.7kg)(9.80m/s^2)(1-\sin31-.10\cos31)/5.4kg=2.0m/s^2\]

OpenStudy (anonymous):

I also know the component forms of box A is Fy=22.68N and Fx=13.62N

OpenStudy (anonymous):

both boxes are in contact?

OpenStudy (anonymous):

Both boxes are connected to a string on a pulley, Box B is hanging from the pulley and box A is on an incline.

OpenStudy (anonymous):

u can draw the free-body diagram showing the forces acting on the box A and B?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

fig isso we have three equation:\[m _{B}g-T=m_{B}a\]\[m_{A}gsin \theta-\mu_{k}m_{A}gcos \theta=m_{A}a\]\[m_{A}+m_{B}=2.7\]so we can solve for mB & mA